0.593 g of magnesium metal was weighed into a styrofoam calorimeter. Next, 60 mL of 1.0 M hydrochloric acid, weighing 60.488 g and having an initial temperature ( i.e. T) of 20.8 oC, was added to the calorimeter. The solution was then stirred and temperatures were recorded as the reaction took place and for more than 300 seconds after the system temperature 'peaked' 

  1. If the final mass of the product in the container is 60.899 g, calculate the average of initial (reactants) and final (product) mass.
1
Expert's answer
2021-01-25T04:22:17-0500

Mg + 2HCl = MgCl2 + H2.

m(total)1 = 0.593 + 60.488 = 61,081 g.

m(total)2= 60.899 g.

m1-m2 = 61.081 - 60.899 = 0,182 g.

m(H2) = 0.182 g, n= 0.182 mol.

0.182 x 24 = 4.368 g. Mg. (4.368+0.182=4,961 g.)

0.182 x 36.5 = 6,643 g.